OpenStudy (charlotteeporte):

Simplify this expression: csc(-x) over 1+tan^2 (x) (Over meaning its a fraction)

OpenStudy (charlotteeporte):

Choices are... tan(x) -cos(x)tan(x) -cos(x)cot(x) sin(x)tan(x)

OpenStudy (dayakar):

1+tan^2 x = sec^2 x [ identity] ------(1) csc (-x) = csc x = 1/sinx -------(2) sec x = 1/cos x -------(3) \[\frac{ \csc (-x) }{ 1 + \tan ^{2}x } = \frac{ \csc x }{ \sec ^{2}x}\] [ from (1) and (3) ]

OpenStudy (dayakar):

\[= \frac{ \frac{ 1 }{ \sin x } }{ \frac{ 1 }{ \cos ^{2}x } }\] [using (2) and (3) \[= \frac{ \cos ^{2}x }{ sinx }\]

OpenStudy (dayakar):

\[=\frac{ \cos x \times cosx }{ \sin x}\] = cotx cos x

OpenStudy (triciaal):